OpenStudy (anonymous):

write y+7=9/10(x+3) in standard form

OpenStudy (anonymous):

multiply by 10 10y+70=9x+27 subtract 10 y 70 = 9x - 10y + 27 subtract 27 43 = 9x - 10y